INFRARED RESONANCE ENHANCED PHOTODISSOCIATION SPECTROSCOPY OF Ni^{+}(CO_{2})_{n}, Ni^{+}(CO_{2})_{n}Ar, AND Ni^{+}(O_{2})(CO_{2})_{n} CLUSTERS

\underline{N.R. WALKER}, G.A. GRIEVES, R.S. WALTERS and M. A. DUNCAN, University of Georgia, Department of Chemistry, Athens, GA, 30602-2556.

Ni^{+}(CO_{2})_{n} and Ni^{+}(CO_{2})_{n}Ar ion-molecule complexes are produced by laser vaporization in a pulsed nozzle source and studied with mass-selected infrared resonance-enhanced photodissociation spectroscopy (IR-REPD). Photofragment yield is measured as a function of energy in the region of the asymmetric stretch vibration of CO_{2} (2349 cm^{−1}). Fragmentation of Ni^{+}(CO_{2})_{n} complexes proceeds by the loss of intact CO_{2} units. The mixed clusters fragment by loss of Ar, permitting higher photofragment yields and sharper spectra than those provided by the corresponding pure Ni^{+}(CO_{2})_{n} complexes. Spectra for the n≤4 complexes depict asymmetric stretches to the blue of free CO_{2} while solvation effects are observed in the larger complexes. An additional blue-shifted band is observed in complexes with n≥6, indicating that the Ni^{+} ion participates in a cluster-assisted chemical reaction producing a nickel oxide. This inference is supported by the results of experiments on Ni^{+}(O_{2})(CO_{2})_{n} complexes, where a blue-shifted band is also observed.